Consider the following data: Nominal GDP for 1993 was $6553 billion, as compared to $6244 for 1992. The GDP deflator for 1993 was 102.6, as compared to 100.0 for 1992. Calculate real GDP for 1992 and 1993, in 1992 prices. Calculate rates of growth of nominal GDP and real GDP for 1993. What was the rate of inflation (as measured by GDP deflator) for 1993?
1
Expert's answer
2020-05-20T09:58:19-0400

"RealGDP=Nominal GDP\/Deflator"

"Real GDP(1992)=\$6244/100.0*100=\$6244"

"Real GDP(1993)=\$6553/102.6*100=\$6387"

"Rate=(Final-Initial)\/Initial"

"Rate (RealGDP)=(\$6387-\$6244)/\$6244*100=2.2" %

"Rate (NominalGDP)=(\$6553-\$6244)/\$6244*100=4.9" %

Rate of inflation for 1993 was 2.6%